# A PNP ion channel deep learning solver with local neural network and finite element input data

### v1

@@ 2.3 The PNPic deep learning solver
where the $i$th grid point $x^{i}=a+ih_{1}$ with the mesh size $h_{1}=(b-a)/(N_{1}+1)$. We refine the grid to obtain another uniform coarse grid with the mesh size $h_{2}=h_{1}/2$. We then solve the two nonlinear finite element systems of Eqs. 4 and 5 derived from these two coarse grids by our PNPic finite element software, respectively, producing two low cost PNPic finite element solutions as the input data for the local neural network scheme.
@@
which, respectively, denote the values of $u$ at the mesh points $x^{i^{\prime}-2},x^{i^{\prime}}$, and $x^{i^{\prime}+2}$ of the second coarse grid, which are at the same locations of the mesh points $x^{i-1},x^{i}$, and $x^{i+1}$ of the first coarse grid. To this end, the input data set $U_{i}$ is obtained in the form
$\begin{array}[]{ll}U_{i}=&\{\phi_{i-1},\phi_{i},\phi_{i+1},c_{1,i-1},c_{1,i},c_{1,i+1},c_{2,i-1},c_{2,i},c_{2,i+1},h_{1},\\ &\;\;\phi_{i^{\prime}-2},\phi_{i^{\prime}},\phi_{i^{\prime}+2},c_{1,i^{\prime}-2},c_{1,i^{\prime}},c_{1,i^{\prime}+2},c_{2,i^{\prime}-2},c_{2,i^{\prime}},c_{2,i^{\prime}+2},h_{2}\},\quad i=1,2,\ldots,N_{1}.\end{array}$ (12)
@@
In the neural network training process, we use the standard loss function defined by
$\mbox{Loss}=\displaystyle\sum_{sample}\sum_{i=1}^{N_{1}}\left\|(\mbox{output of the deep learning solver at $U_{i}$})-(\mbox{reference solution at $x^{i}$})\right\|^{2},$ (13)
@@
In the training process, we also use multiple steps of Adams optimizer and a final step of the limited-memory BFGS (L-BFGS) optimization algorithm for minimizing the loss function to determine the weight and bias parameters of the neural network optimally. By default, we set the learning rate of L-BFGS as $10^{-5}$ and the learning rate of Adams as $10^{-4}$. The neural network of our PNPic solver is implemented using Pytorch 2.1.0 on Google Colab.
In summary, our PNPic deep learning solver combines the PNPic finite element solver with the local neural network scheme. Its input data sets come from the two coarse grid PNPic finite element solutions and it outputs a predicted PNPic solution in very high accuracy. Moreover, it can be used to predicted a PNPic solution in high accuracy for a family of PNPic model problems provided that it has been trained by using various PNPic solution data generated by using perturbed parameters, perturbed subdomains, and perturbed boundary values, etc.
@@ 2.4 Numerical test results
In numerical tests, we set the interval $[a,b]$ with $a=-5$ and $b=8.5$. We then set the interface numbers $x_{1}=0$, $x_{2}=0.2$, $x_{3}=1.3$, $x_{4}=2.3$, and $x_{5}=3.5$. For clarity, we list default parameter values and coefficient functions of the PNPic model in Table 1. From this table, we can get that the radius function $r(x)$ is a constant function with value 0.5 in $0\leq x\leq 3.5$, a linear function in the interior bath, defined by the two values $r(-5)=5.5$ and $r(0)=0.5$, and another linear function determined by $r(3.5)=0.5$ and $r(8.5)=5.5$ in the exterior bath. These intervals and parameter values come from a study of the ${}^{+}$ channel [19, 9]. We construct two coarse grid meshes with the mesh sizes $h_{1}=1/40$ and $h_{2}=1/80$ and one fine grid mesh with the mesh size $h_{ref}=1/640$. These meshes include the interface points $x_{i}$ for $i=1,2,3,4,5$ as mesh grid points. We solve the PNPic model 2.1 by our software package based on these three meshes.
In these tests, we set the charge density function $\rho$ within each subdomain of the ion channel pore in the expression
@@
[…]
@@
where ${\cal T}$ denotes a set of perturbation cases done in the prediction/validation tests, $U^{i}_{p}$ denotes a solution of the PNPic model predicted by our PNPic deep learning solver in the $i$th perturbation, $U^{i}_{h2}$ is a numerical solution of the PNPic model on the second coarse grid mesh for training in the $i$th perturbation, and $U^{i}_{ref}$ is the numerical solution on the finest grid for the $i$th perturbation. The smaller the $\tau_{i}$ value, the more improvement is done by our PNPic deep learning solver in comparison with the input solution data produced by our PNPic finite element solver. Our choice of the norm in the reduction factor is the $l_{2}$-norm which is consistent with the loss function 13 used in training.
Table 1: The subdomain intervals of a ${}^{+}$ channel pore and the model parameters used in database generation.
@@
In the first type of tests, we perturb only one parameter, say $p$, of the PNPic model from $p$ to $\hat{p}$ with $\hat{p}=(1+\lambda)p$ for $\lambda=0,\pm 0.01,\pm 0.02,\ldots,\pm 0.09,\pm 0.1$ while retaining all the other parameters. That is, a total of $21$ tests are carried out and in each test we generate three different PNPic finite element solutions based on the two coarse grid meshes and the fine grid mesh. We then split the generated data set into the training data set that includes the solution data from the tests using $\hat{p}=(1+\lambda)p$ for $i=0,\pm 0.1$ and the testing data set that includes the solution data from the tests using $\hat{p}=(1+\lambda)p$ for $\lambda=\pm 0.01,\pm 0.02,\ldots,\pm 0.08,\pm 0.09$. In our numerical tests, we set $p=D_{c},D_{f},\epsilon_{c},\epsilon_{f},Q_{b}$ that represent the default values from Table 1
$D_{c}=0.4,\quad D_{f}=0.4,\quad\epsilon_{c}=30,\quad\epsilon_{f}=30,\quad Q_{b}=-4.$
The test results are reported in Figures 2 to 3. We employed 5 neural networks for these five single parameter perturbation cases, respectively, and obtained the reduction factor of no larger than $\approx 0.17$ in all these five testing cases as shown in Figure 2.
In the second type of tests, we perturb two parameter values simultaneously. Let $p_{1}$ and $p_{2}$ denote those two parameters to be disturbed from their default values in Table  1. We then disturb them to $(1+\delta_{1})p_{1}$ and $(1+\delta_{2})p_{2}$ for $\delta_{1}=0,\pm 0.1$ and $\delta_{2}=0,\pm 0.1$. The two parameters selected for tests are the two permittivity constants in the cavity and filter regions, $\epsilon_{c}$ and $\epsilon_{f}$, the two diffusivity constants in the cavity and filter regions, $D_{c}$ and $D_{f}$, and the two boundary values of the two concentrations $c_{1}$ and $c_{2}$, $c_{i,a}$ and $c_{i,b}$ with $a=-5$ and $b=8.5$ for $i=1,2$. As done in the case of single parameter case, we dedicate one neural network to each choice of two parameter values, hence the total of 4 neural networks. The specific testing cases and reduction factors are reported in Tables 2 and 3.
@@
Figure 2: Numerical solution accuracy improvements by our PNPic deep learning solver in terms of the reduction factor $\tau_{i}$ defined in Eq.14. Here 18 prediction test cases were done by our PNPic deep learning solver for the five parameters $D_{c},D_{f}$, $\epsilon_{c}$, $\epsilon_{f}$, and $Q$, respectively.
@@
In our final numerical test, we illustrate that our PNPic deep learning solver can also accurately predict solutions corresponding to shifted interface locations. To this end, we adjusted the interface number $x_{5}=3.5$, which separates the filter and exterior bath regions, by adding the perturbations ±$\pm 4h_{1},\pm 3h_{1},\pm 2h_{1},\pm h_{1}$. Our training data set consists of the cases using $-4h_{1},-h_{1},2h_{1},4h_{1}$. In contrast with the previous experiments, we decreased the training gap (relative difference between adjacent two parameters in reference solutions used for training) from $10\%$ to approximately $5\%$ due to solution sensitivity on the interface location. Our particular selection of the training data set is guided by a comparison of the coarse grid input solutions to take into account the various effects of perturbing the interface location on the solutions. The results for the testing cases (including the case for the unperturbed interface location) are reported in Table 4.
@@
Table 4: Accuracy improvement of our PNPic deep learning solver for testing cases with the interface located at $x_{5}+p$ for $p=-3h_{1},-2h_{1},0,h_{1},3h_{1}$. Here $\tau_{i}$ is the reduction factor of Case $i$ defined in Eq.14.
@@ 3 Discussion
We can see from Figure 2 that our PNPic deep learning solver can produce more accurate numerical solution of the PNPic model than the input solution computed by the PNPic finite element solver. Given that the underlying finite element solver is the second order accurate, the reduction factor of $0.17$ amounts to greater accuracy than what would be expected if we were to reduce the finer mesh size from $h_{2}$ to $h_{2}/2$. The closer the parameter values are to the default ones from 1, the smaller the reduction factors, (larger accuracy improvement) which can potentially be attributed to the symmetry of the training data sets.
@@
Tables 2 and 3 illustrate that the reduction factors for the testing cases of perturbing two parameters are comparable to the single parameter cases. The training cost becomes more expensive as the number of training data set increases with the number of parameter values. However, the training gaps are around $5\%$ for different interface positions used in training and no smaller than $10\%$ in other parameters set for training. This allows us to use about tens of fine reference solutions corresponding to different parameter values and domains to cover a relatively large parameter space ($O(1)$ in size for the convex hull of the parameters used in training.) The trained neural network is able to predict the high-fidelity solution for any parameter values in the parameter space.
@@ 4 Conclusion
We have developed a PNPic deep learning solver, which is defined as a combination of a new 1D PNPic finite element solver with a local neural network, and is the first to adapt NNLCI to a different PDE system with convection-diffusion equations and elliptic equations. In addition, it is the first time for us to use NNLCI to deal with multiple subdomains connected with complicated interface conditions. Instead of assigning a neural network to each subdomain and addressing the interface conditions between these neural networks, we only use one single local neural network with no need to deal with the interface conditions in the neural network. The simplicity of the methodology provides us with great robustness and efficiency for practical use and implies a breakthrough for this type of complicated PDE problem.
With the success of this work, we will develop a 3D PNP ion channel deep learning solver for solving a 3D PNP ion channel model and train it by using the data reflecting various ion channel protein molecular structures. We expect that the 3D solver will have orders of magnitude reduction in complexity and training costs in comparison to the conventional global neural network approach. For example, if the number of coarse-grid cells is about one-eighth of that of the reference-grid cells as in the paper, the number of coarse-grid cells will be 1/64 in 2D and 1/512 in 3D of those of the reference-grid cells. This implies about two and three orders of magnitude reduction in complexity in 2D and 3D, respectively.
Author Contributions
Author 1 conducted training, prediction, and some data generation. Author 2 carried out all simulations and some data generation, and analyzed the data. Authors 1-5 contributed to the methodology and writing of the article.
@@ Acknowledgments
We validated our PNPic software package by using an analytical solution of the following PNP system
$-A(x)^{-1}\frac{d}{dx}\left[\epsilon(x)A(x)\frac{d\phi(x)}{dx}\right]=10\left[z_{1}c_{1}(x)+z_{2}c_{2}(x)+g(x)\right],$ $0<x<1,$
[…]
$\phi(0)=0,\quad\phi(1)=-1,\quad c_{1}(0)=0,\quad c_{1}(1)=0,\quad c_{2}(0)=0,$ $c_{2}(1)=0,$
@@
It is easy to validate that the analytical solution of System Appendix: Validation tests by an analytical solution is given by
@@
Therefore, this system can be used to validate our PNPic finite element software package.
We did so by solving the system via our package based on a uniform mesh of the interval $[0,1]$ with the mesh size $h=1/20,1/40,1/80,1/160$, and 1/320, respectively. We then calculated an absolute error, $E(u_{h},u)$, between the finite element solution $u_{h}$ and analytical solution $u$ for $u=c_{1},c_{2},$ and $\phi$ by the expression
@@
From finite element theory, it is known that a linear finite element method has a quadratic order of convergence rate (i.e. $p=2$) when we calculate the error by Eq.17, which can be approached, monotonically, by a sequence of convergence order estimations, $\{p_{u,h_{k}}\}$, such that $p_{u,h_{k}}\rightarrow 2$ as $k\rightarrow\infty$. Here $p_{u,h_{k}}$ is defined by
@@
and $h_{k}$ denotes the mesh size of the $k$th grid mesh satisfying
$h_{1}>h_{2}>\ldots>h_{k-1}>h_{k}>\ldots>0.$
We calculated $E(u_{h_{k}},u)$ and $p_{u,h_{k}}$ for $u=c_{1},c_{2},\phi$ and $k=1$ to 5 and listed the results in Table 5. From Table 5 it can be seen that $E(u_{h_{k}},u)$ approaches zero and $p_{u,h_{k}}$ approaches 2, monotonically, for $u=c_{1},c_{2},$ and $\phi$ as the grid size $h_{k}$ decreases from $1/20$ to $1/320$. These numerical results coincide with finite element theory, validating our PNPic software package.
Table 5: Test results produced by our PNPic finite element software package from solving System Appendix: Validation tests by an analytical solution.

### v2

@@ 2.3 The PNPic deep learning solver
where the $i$th grid point $x^{i}=a+ih_{1}$ with the mesh size $h_{1}=(b-a)/(N_{1}+1)$. We refine the grid to obtain another uniform coarse grid with the mesh size $h_{2}=h_{1}/2$. We then solve the two nonlinear finite element systems of equations  (4) and (5) derived from these two coarse grids by our PNPic finite element software, respectively, producing two low cost PNPic finite element solutions as the input data for the local neural network scheme.
@@
which, respectively, denote the values of $u$ at the mesh points $x^{i^{\prime}-2},x^{i^{\prime}}$, and $x^{i^{\prime}+2}$ of the second coarse grid at the same locations of the mesh points $x^{i-1},x^{i}$, and $x^{i+1}$ of the first coarse grid. To this end, the input data set $U_{i}$ is obtained in the form
$\begin{array}[]{ll}U_{i}=&\{\phi_{i-1},\phi_{i},\phi_{i+1},c_{1,i-1},c_{1,i},c_{1,i+1},c_{2,i-1},c_{2,i},c_{2,i+1},h_{1},\\ &\;\;\phi_{i^{\prime}-2},\phi_{i^{\prime}},\phi_{i^{\prime}+2},c_{1,i^{\prime}-2},c_{1,i^{\prime}},c_{1,i^{\prime}+2},c_{2,i^{\prime}-2},c_{2,i^{\prime}},c_{2,i^{\prime}+2},h_{2}\}.\end{array}$ (13)
@@
In the neural network training process, we define the loss function by
@@
$-({\text{reference solution at }}x^{i})\|^{2},$
[…]
In the training process, we start with multiple steps of Adams optimizer and then switch to the limited-memory BFGS (L-BFGS) optimization algorithm for minimizing the loss function to determine the weight and bias parameters of the neural network optimally. By default, we set the learning rates as $10^{-4}$ and $10^{-5}$ for Adams and L-BFGS, respectively. The neural network of our PNPic deep learning solver is implemented on Google Colab using Pytorch 2.1.0.
In summary, our PNPic deep learning solver is a novel combination of our efficient PNPic finite element solver with our effective local neural network scheme. Its input data sets come from the two PNPic coarse grid finite element solutions and it outputs a predicted PNPic solution in very high accuracy. Compared with the traditional finite element method, which can only find one approximate solution of the model, our PNPic deep learning solver can predict a family of PNPic solutions with high accuracy for a family of PNPic model problems provided that it is well trained by using the PNPic solution data generated from perturbed parameters, perturbed subdomains, and perturbed boundary values, etc.
@@ 2.4 Numerical test results
To demonstrate that our PNPic deep learning solver can be used to produce an approximate solution of the PNPic model in high accuracy for a family of model problems, we did different numerical tests in this section. Following [9], we set the interval $[a,b]$ with $a=-5$ and $b=8.5$. We then set the interface numbers $x_{1}=0$, $x_{2}=0.2$, $x_{3}=1.3$, $x_{4}=2.3$, and $x_{5}=3.5$. For clarity, we list default parameter values and coefficient functions of the PNPic model in Table 1. From this table, we can get that the radius function $r(x)$ is a constant function with value 0.5 in $0\leq x\leq 3.5$, a linear function in the interior bath, defined by the two values $r(-5)=5.5$ and $r(0)=0.5$, and another linear function determined by $r(3.5)=0.5$ and $r(8.5)=5.5$ in the exterior bath. These intervals and parameter values come from a study of the K+ channel [9, 4]. We construct two coarse grid meshes with the mesh sizes $h_{1}=1/40$ and $h_{2}=1/80$ and one fine grid mesh with the mesh size $h_{ref}=1/640$. These meshes include the interface points $x_{i}$ for $i=1,2,3,4,5$ as mesh grid points. We solve the PNPic model (1) by our software package based on these three meshes.
Following [9], we set the charge density function $\rho$ within each subdomain of the ion channel pore in the expression
@@
where $Q$, $\ell$, and $A(x)$ denote the charge amount (in Coulomb) within the subdomain, the subdomain length, and the subdomain cross-sectional area function, respectively.
@@
where ${\cal T}$ denotes a set of perturbation cases done in the prediction/validation tests, $U^{i}_{p}$ denotes a solution of the PNPic model predicted by our PNPic deep learning solver in the $i$th perturbation, $U^{i}_{h2}$ is a numerical solution of the PNPic model on the second coarse grid mesh for training in the $i$th perturbation, and $U^{i}_{ref}$ is the numerical solution on the finest grid for the $i$th perturbation. The smaller the $\tau_{i}$ value, the more improvement is done by our PNPic deep learning solver in comparison with the input solution data produced by our PNPic finite element solver. Our choice of the norm in the reduction factor is the $l_{2}$-norm which is consistent with the loss function (14) used in training.
Table 1: The subdomain intervals of a K+ channel pore and the model parameters used in database generation.
@@
In the first type of tests, we perturb only one parameter, say $p$, of the PNPic model from $p$ to $\hat{p}$ with $\hat{p}=(1+\lambda)p$ for $\lambda=0,\pm 0.01,\pm 0.02,\ldots,\pm 0.09,\pm 0.1$ while retaining all the other parameters. That is, a total of $21$ tests are carried out and in each test we generate three different PNPic finite element solutions based on the two coarse grid meshes and the fine grid mesh. We then split the generated data set into the training data set that includes the solution data from the tests using $\hat{p}=(1+\lambda)p$ for $i=0,\pm 0.1$ and the testing data set that includes the solution data from the tests using $\hat{p}=(1+\lambda)p$ for $\lambda=\pm 0.01,\pm 0.02,\ldots,\pm 0.08,\pm 0.09$. In our numerical tests, we set $p$ as a value of $D_{c},D_{f},\epsilon_{c},\epsilon_{f}$, or $Q$ as follows:
$D_{c}=0.4,\quad D_{f}=0.4,\quad\epsilon_{c}=30,\quad\epsilon_{f}=30,\quad Q=-4e_{c},$
where $Q$ is the total charge amount in the buffer region. In these tests, we used five different neural networks for the five single parameter perturbation cases. We calculated the reduction factor $\tau_{i}$ of (15) for each test case and reported these values in Figure 2. From this figure we can see that all the values of the reduction factor $\tau_{i}$ were less than $0.17$, indicating that our predictions have high accuracy.
In the second type of test, we perturb two parameter values simultaneously. Let $p_{1}$ and $p_{2}$ denote those two parameters to be disturbed from their default values in Table 1. We then disturb them to $(1+\delta_{1})p_{1}$ and $(1+\delta_{2})p_{2}$ for $\delta_{1}=0,\pm 0.1$ and $\delta_{2}=0,\pm 0.1$. The two parameters selected for tests are the two permittivity constants in the cavity and filter regions, $\epsilon_{c}$ and $\epsilon_{f}$, the two diffusivity constants in the cavity and filter regions, $D_{c}$ and $D_{f}$, and the two boundary values of the two concentrations $c_{1}$ and $c_{2}$, $c_{i,a}$ and $c_{i,b}$ with $a=-5$ and $b=8.5$ for $i=1,2$. As done in the case of single parameter case, we dedicate one neural network to each choice of two parameter values, hence the total of 4 neural networks. The specific testing cases and reduction factors are reported in Tables 2 and 3.
@@
Figure 2: Numerical solution accuracy improvements by our PNPic deep learning solver in terms of the reduction factor $\tau_{i}$ defined in (15). Here 18 prediction test cases were done by our PNPic deep learning solver for the five parameters $D_{c},D_{f}$, $\epsilon_{c}$, $\epsilon_{f}$, and $Q$, respectively.
@@
In our final numerical test, we illustrate that our PNPic deep learning solver can also accurately predict solutions corresponding to shifted interface locations. To this end, we adjusted the interface number $x_{5}=3.5$, which separates the filter and exterior bath regions, by adding the perturbations ±$\pm 4h_{1},\pm 3h_{1},\pm 2h_{1},\pm h_{1}$. Our training data set consists of the cases using $-4h_{1},-h_{1},2h_{1},4h_{1}$. In contrast with the previous experiments, we decreased the training gap (relative difference between adjacent two parameters in reference solutions used for training) from $10\%$ to approximately $5\%$ due to solution sensitivity on the interface location. Our particular selection of the training data set is guided by comparison of the coarse grid input solutions to take into account the various effects of perturbing the interface location on the solutions. The results for the testing cases (including the case for the unperturbed interface location) are reported in Table 4.
@@
Table 4: Accuracy improvement of our PNPic deep learning solver for testing cases with the interface located at $x_{5}+p$ for $p=-3h_{1},-2h_{1},0,h_{1},3h_{1}$. Here $\tau_{i}$ is the reduction factor of Case $i$ defined in (15).
@@ 3 Discussion
We can see from Figure 2 that our PNPic deep learning solver can produce a more accurate numerical solution of the PNPic model than the input solution computed by the PNPic finite element solver. Given that the underlying finite element solver is second order accurate, the reduction factor of $0.17$ amounts to greater accuracy than what would be expected if we were to reduce the finer mesh size from $h_{2}$ to $h_{2}/2$. The closer the parameter values are to the default ones from Table 1, the smaller the reduction factors, (larger accuracy improvement) which can potentially be attributed to the symmetry of the training data sets.
@@
Tables 2 and 3 illustrate that the reduction factors for the testing cases of perturbing two parameters are comparable to the single parameter cases. The training cost becomes more expensive as the number of training data sets increases with the number of parameter values. However, the training gaps are around $5\%$ for different interface positions used in training and no smaller than $10\%$ for other parameters set for training. This allows us to use tens of fine reference solutions corresponding to different parameter values and domains to cover a relatively large parameter space ($O(1)$ in size for the convex hull of the parameters used in training.) The trained neural network can predict the high-fidelity solution for any parameter values in the parameter space.
@@ 4 Conclusion
We have developed a PNPic deep learning solver, which is defined as a combination of a new 1D PNPic finite element solver with a local neural network, and is the first to adapt NNLCI to a different PDE system with convection-diffusion equations and elliptic equations. In addition, it is the first time for us to use NNLCI to deal with multiple subdomains connected with complicated interface conditions. Instead of assigning a neural network to each subdomain and addressing the interface conditions between these neural networks, we only use one single local neural network with no need to deal with the interface conditions in the neural network. The simplicity of the methodology provides us with great robustness and efficiency for practical use, and implies a breakthrough for this type of complicated PDE problem.
NNLCI was originally introduced for predicting weak solutions of 1D Euler systems containing shocks, contacts and their interactions in [11]. This 1D work paves the way for the successful application of NNLCI to the 2D Euler systems in [12]. The extension of NNLCI from 1D to a higher dimension is completely natural with no change in methodology. The most critical part of the work is to first demonstrate that NNLCI works for completely different types of equations in a multi-domain environment in 1D. Based on the prior works mentioned above, we have strong reasons to believe that the highly successful approach developed for the 1D PNPic deep learning solver also works for the development of a 3D PNPic deep learning method for solving a 3D PNPic model. We will develop a 3D PNP ion channel deep learning solver for solving a 3D PNP ion channel model and train it by using the data reflecting various ion channel protein molecular structures. We expect that the 3D solver will not only capture additional spatial details and phenomena but also have orders of magnitude reduction in complexity and training costs in comparison to fine-grid simulations and conventional global neural network approaches respectively. For example, if the coarse-grid cell size is at least $8$ times as large as that of the reference-grid as in the paper, the number of coarse-grid cells will be 1/64 in 2D and 1/512 in 3D of those of the reference-grid cells. This implies about two and three orders of magnitude reduction in complexity in 2D and 3D, respectively. The choice between 1D and 3D PNPic deep learning methods depends on the specific research questions, available computational resources, and required simulation accuracy levels in particular study.
@@ Acknowledgments
Author Contributions
Lee conducted neural network training and analyzed simulation results. Chao and Lee generated data for training and prediction. Lee, Chao, Cobb, and Xie contributed to the package development. Authors 1 to 5 contributed to the manuscript writing and editing. Liu and Xie designed methodology and managed the research activities.
@@
We validated our PNPic software package by using an analytical solution of the following PNP ion channel model:
$-A(x)^{-1}\frac{d}{dx}\left[\epsilon(x)A(x)\frac{d\phi(x)}{dx}\right]=10\left[z_{1}c_{1}(x)+z_{2}c_{2}(x)+g(x)\right],\quad 0<x<1,$
$A(x)^{-1}\frac{d}{dx}\left[A(x)D_{i}(x)\left(\frac{dc_{i}(x)}{dx}+40z_{i}c_{i}(x)\frac{d\phi(x)}{dx}\right)\right]=f_{i}(x),\quad i=1,2,\quad 0<x<1,$
$\phi(0)=0,\quad\phi(1)=-1,\quad c_{1}(0)=0,\quad c_{1}(1)=0,\quad c_{2}(0)=0,\quad c_{2}(1)=0,$
@@
It is easy to find the analytical solution of the above system as
@@
In this validation test, we solved this system based on five uniform meshes of the interval $[0,1]$ with the mesh size $h=1/20,1/40,1/80,1/160$, and $1/320$, respectively. We then calculated an absolute error, $E(u_{h},u)$, between the finite element solution $u_{h}$ and analytical solution $u$ for $u=c_{1},c_{2},$ and $\phi$ by the expression
@@
From the finite element theory, it is known that a linear finite element method has a quadratic order of convergence rate (i.e. $O(h^{p})$ with $p=2$). The order $p$ can be approached monotonically by a sequence of estimations, $\{p_{u,h_{k}}\}$, defined by
@@
where $h_{k}$ denotes the mesh size of the $k$th grid mesh satisfying
$h_{1}>h_{2}>\ldots>h_{k-1}>h_{k}>\ldots>0,$
and $u_{h_{k}}$ denotes the solution defined on the $k$th mesh. We calculated $E(u_{h_{k}},u)$ and $p_{u,h_{k}}$ for $u=c_{1},c_{2},\phi$ using $h_{k}=1/(2^{k}10)$ for $k=1$ to 5 and listed the results in Table A.
From Table A, we can see that $E(u_{h_{k}},u)$ approaches zero and $p_{u,h_{k}}$ approaches 2, monotonically, for $u=c_{1},c_{2},$ and $\phi$ as the grid size $h_{k}$ decreases from $1/20$ to $1/320$. These numerical results coincide with the finite element theory, partially validating our PNPic software package.
@@
Table A: Validation test results produced by our PNPic software package.